OpenStudy (anonymous):

Use y=mx+b form to write the equation of a line with a slope -3 and y intercept 0.

OpenStudy (mathstudent55):

slope = m = -3 y-intercept = b = 0 Just insert those values in y = mx + b and simplify.

OpenStudy (anonymous):

Whats X?

OpenStudy (mathstudent55):

y ans x are part of the equation and stay there.

OpenStudy (anonymous):

Can you explain it better please?

OpenStudy (anonymous):

x is the value of x in the ordered pair (x,y). Slope intercept form is the relationship between X and y.

OpenStudy (mathstudent55):

y = mx + b with m = -3 and b = 0 y = (-3)x + 0 y = -3x

OpenStudy (anonymous):

So saying y=3x means that for any x value, the corresponding y value is 3times that x value.

OpenStudy (anonymous):

Oh okay